# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 24.2 Lower bounds on the bit operation complexity
Next strengthen the lower bound for Problem 0.
Observation 24.2.
Let $p(x)=(x-v_{1})^{j}f(x)$ for an integer $j>0$ and a monic polynomial $f(x)$ of degree $d-j$. Fix a real $b$. Then the polynomial $p_{j}(x)=p(x)-2^{-jb}f(x)$ has the $j$ zeros $v_{1}+\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$.
Observation 24.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to approximate within $2^{-b}$ a $j$-multiple zero of $p$ or its zero in a cluster of $j$ zeros, and so a root-finder must process at least $jb$ bits of the coefficient $p_{d-j}$. Sum these bounds for $j=1,\dots,d$ and obtain
@@
One must access at least $0.5(d+1)db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(d+1)db$ bit operations to approximate even its single within $2^{-b}$.
The lower bounds of this subsection imply that the algorithm of [96, 103] and its extension supporting Thm. 24.1 solve Problems 0 and F of large size by using optimal number of bit operations up to polylogarithmic factors.
24.3 Linking Factorization and Root-finding (with more details)
Next we prove Thm. 24.2, by applying the results of [134, Cor. 4.3 and Sec. 19].
@@ 24.3 Linking Factorization and Root-finding (with more details)
Based on [134, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
@@
[136, Thm. 2.7], which we reproduce in Sec. 25.4 as Thm. 25.3, shows an alternative way to relaxing both assumptions – that $r_{1}\leq 1$ and $p(x)$ is monic.
Schönhage in [134, Sec. 19] has greatly strengthened the estimate of Thm. 24.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [134, Eqn. (19.8)]) the following result.
@@
By extending these estimates we strengthen Thm. 24.4 as follows.
@@
Under bounding cluster sizes of Def. 24.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$. Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{1}(b,d)\leq\mathbb{B}_{2}(\tilde{O}(b+d),d)$.
For converse reduction of Problem F to Problem 0 recall [134, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
@@
Now consider the products $p=gh$ for $g=x-z_{j}$, for $j=1,\dots,d$. For a fixed $j$, an error of at most $\epsilon/2^{d}$ for the root $z_{j}$ contributes at most $\epsilon||h||_{1}$ to overall factorization error of Problem F. Therefore, $||h||_{1}\leq 2^{d-1}||p||_{1}/||x-z_{j}||_{1}\leq 2^{d-1}||p||_{1}$ by virtue of Thm. 24.5. Hence the approximation errors of all the $d$ roots together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
@@
If $b=O(b-d-\log(d))$, then Problem F for an error bound $1/2^{b}$ can be reduced at a dominated bit operation cost to Problem 0 for an error bound $1/2^{\widehat{b}}$ and a sufficiently large $\widehat{b}=O(b+d+\log(d))$, and so $\mathbb{B}_{F}(b,d)=O(\mathbb{B}_{0}(b,d))$.
It follows that the complexity estimate of Thm. 1.3 can be also applied to the solution of Problem F of factorization provided that $b=O(b-d-\log(d))$. Then again, the bit operation complexity of this solution of Problem F is optimal (up to polylogarithmic factors) under the random root model provided that $d=\tilde{O}(b)$ (cf. Observation 24.1).
24.4 Two root-finders of 2016 and 2018
Becker et al in [23, 24] claimed that their extension of subdivision root-finders of [128, 101] (i) runs within the same near optimal bit operation bounds of [96, 103] but (ii) unlike [96, 103] is easy to implement. This seemed to diminish incentive for further research in this area, but were the claims valid?
Implementation of the root-finder of [23] by Imbach et al in [64] provided no empirical support to optimality claimed by Becker et al because the root-finder of [63], using some novel techniques from [112] (see [63, Sec. 1.2]), runs much faster. Unfortunately, formal support for optimality claimed by Becker et al has been missing as well.
Indeed, for all proofs the paper [23] refers to a preprint of [24], with a distinct list of the authors and with the main algorithms and the Main Theorem significantly different from those of [23].
Furthermore, as we pointed out in the introduction, the the bit-complexity specified in the Main Theorem of [24] has order of $mbd^{2}$, that is, exceeds the near optimal bound of [96, 103] by a factor of $d$.
Can one salvage the claim by Becker et al in another way? Next we argue that this can be problematic because the algorithms of [23, 24] for Problem 1m are some variants of subdivision iterations. The basic tool and the main algorithmic novelty of the variant of subdivision in [96, 103] are e/i tests based on the following theorem by Pellet [88], which extended the previous results by Cauchy and Rouché:
Theorem 24.6.
Given a polynomial $(x)$ of (1.1) with complex coefficients and $p_{g}=1$. Let the polynomial $s(x)=x^{d}+p_{d-1}x^{d-1}+\cdots+|p_{i+1}|x^{i+1}-|p_{i}|x^{i}|+p_{i-1}|x^{i-1}+\cdots+|p_{0}|$ have two distinct positive roots $r$ and $R$, $r<R$. Then the polynomial $p$ has exactly $i$ zeros in or on the circle $C(0,r)=\{z:~{}|z|=r\}$ and no zeros in the annulus $A(0,r,R)=\{z:~{}r<|z|<R\}$.
[23, 24] apply Pellet’s theorem on each subdivision step to more than $4m$ polynomials $t(x)=\sum_{i=0}^{d}t_{i}x^{i}$ obtained from the polynomial $p$ by means of shifting the variable $x$, its scaling by up to $2^{b}$, and root-squaring.
To apply Pellet’s theorem to $t(x)$ one must compute the values $|\frac{t_{i}}{t_{d}}|$ for $i=0,1,\dots,d-1$. Hence, by virtue of Observation 24.2, one must process at least $0.5(d+1)db$ bits of the coefficients of $t(x)$ to support approximation of its zeros within $1/2^{b}$.
The algorithms of [24] cannot avoid this because they apply e/i tests to the polynomials obtained in $\log(\log(d))$ steps of root-squaring (cf. Sec. 13.3) from polynomials $t_{c,\rho}(x)=p(\frac{x-c}{\rho})$ where $\rho>0$ can be as small as $1/2^{b}$.
Already a single subdivision iteration can perform more than $4m$ soft e/i tests; this means that one must process more than $2(d+1)dmb$ bits of memory and $(d+1)dmb$ bit-operations to solve Problem 1m by applying such a Pellet-based root-finder.
25 Recursive Factorization; extensions to Root-finding and Root Isolation
@@ 25.1 Auxiliary norm bounds
[136, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma 25.3. From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta:=\zeta_{q}:=\exp(\frac{2\pi{\bf i}}{q})$ denoting a primitive $q$-th root of unity, one can see some limitations on strengthening this bound further.
@@ 25.2 Errors and complexity
Next, by following [134, Sec. 5], we estimate the norm of the residual polynomial
@@
Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (25.3)-(25.5) and Lemmas 25.1 and 25.3 and obtain (25.6). ∎
Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (25.6) turns into (25.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (25.4), apply Thm. 25.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (25.1) with
@@
The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (25.7) ensures (25.5), and then (25.6) follows by virtue of Thm. 25.1.
@@
25.3 Overall complexity of Recursive Factorization and extension to Root Isolation
The overall complexity of recursive factorization is dominated by the sum of the bounds on the complexity of all splittings into pairs of factors. The algorithm of [103] adopts recursive factorization of [134] but decreases the overall complexity bound by roughly a factor of $d$. This is due to additional techniques, which ensure that in every recursive splitting of a polynomial into the product of two factors the ratio of their degrees is at most 11. Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [103], and so the overall complexity of root-finding is proportional to that of the first splitting. In contrast, the degrees of a factor in any splitting
in [134] can have low, e.g., 1, and so there can be allows up to $d-1$ splittings overall, with each of the first $0.5d$ of them almost as expensive as the very first one.
@@ 25.4 From factors to roots
By virtue of Thm. 25.3 for $b^{\prime}=O(bd)$ we can bound the bit operation complexity of the solution of Problem 0 by increasing the estimate for the complexity of factorization in Sec. 25.3 by a factor of $d$.
@@
The study in this section covers the solution of Problem 1 for the worst case input polynomial $p$. If its roots are isolated from each other, then the upper estimates of Thm. 25.3 and Cor. 25.2 can be decreased by a factor of $d$ (cf. [103]).
@@ 25.5 Polynomial Root Isolation
Based on his results recalled in this section, Schönhage in [134, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
Problem ISOL. Polynomial Root Isolation. Given a polynomial $p$ of (1.1) that has integer coefficients and is square-free, that is, has only simple roots, compute $d$ disjoint discs on the complex plane, each containing exactly one root.
@@
Our Problem 0 can be viewed as the Problem of Polynomial Root Cluster Isolation. It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem 0 as well as the Algorithms of [96, 103] solve Problem ISOL within a record complexity bounds.
In [134, Sec. 20] Schönhage reduces to factorization Problem F the root-isolation problem of computing $d$ disjoint discs each covering a single simple root of a square-free polynomial with integer coefficients. Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [77, 78, 79]), and Schönhage has estimated that solution of Problem F (e.g., by means of the algorithms of [134, 94, 103]) implies isolation provided that
@@ 26.1 Contents
In the next subsection we recall Newton’s, Weierstrass’s, and Ehrlich’s iterations, comment on their convergence, and acceleration by means of applying fast multipoint evaluation, which is dramatic but involves the coefficients of $p$. In Sec. 26.3 we cover mre limited acceleration of these iterations by means of incorporating FMM, which can be applied to a black box polynomial. In Secs. 26.5 and 27) we cover initialization of Newton’s, Weierstrass’s, and Ehrlich’s iterations, required for their fast global empirical convergence. For Newton’s iterations we only need to compute a disc that contains all zeros of a polynomial $p$ (cf. [71, 58, 143, 130, 141]). This is given by the well-known estimates via the coefficients, but we present simple but novel recipes in the case of black box polynomial $p$. The customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations but is still fast for polynomials represented with their coefficients, as we recall in Sec. 27), where we also point out how this initialization can be extended to fast real and nearly real root-finding. It is still a challenge to extend such an initialization to the case of a black box polynomial.
@@ 26.2 Definitions
1. Definitions. Consider Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s) iterations [71, 58, 130, 143, 148, 31, 67, 38, 3, 22]):
@@
In contrast, Weierstrass’s and Ehrlich’s iterations begin at $d$ points for a $d$th degree polynomial and approximate a subset of its $d$ roots about as fast and almost as slow as they approximate all its $d$ roots.
Near the roots Newton’s and Weierstrass’s iterations converge with quadratic rate, Ehrlich’s iterations with cubic rate [145], but empirically all of them consistently converge very fast globally, right from the start, assuming a proper initialization, which involves the coefficients of $p$ (see [15, 22]). [^37] Formal support for such global convergence is much weaker for Newton’s iterations (see [71, 58, 143], and the references therein) and is completely missing for Weierstrass’s and Ehrlich’s iterations.
All expressions in (26.3) for the functions $N(x)$, $W(x)$, and $E_{j}(x)$ are already presented or can be rewritten as ratios $\frac{u(x)}{v(x)}$ for pairs of polynomials $u(x)$ and $v(x)$, whose coefficients are given or readily available at low cost if the polynomial $p(x)$ is gien with its coeffcients (see [102, Problems 3.6.1 and 3.6.3]). Then multipoint evaluation of such a function amounts to multipoint evaluation of $u(x)$ and $v(x)$ and division of the computed values, and one can accelerate Newton’s, Weierstrass’s, and Ehrlich’s iterations dramatically by incorporating the algorithms of [81, 59].

### v37

@@ 22.2 Linking Factorization and Root-finding: lower bounds on the bit operation complexity
Next we deduce a stronger lower bound for Problem 0.
Observation 22.2.
Let $p(x)=x^{d}+p_{d-j}x^{d-j}$ for an integer $j$, $1<j<d$, and complex $p_{d-j}$ such that $|p_{d-j}|=1/2^{bj}$. Then the polynomial $p(x)$ has the $j$ zeros $\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$, which vanish in the transition to $x^{d}$.
Observation 22.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to distinguish the above $j$ zeros of $p(x)$ from the value 0 or from the zeros of the cluster of $j$ zeros of the polynomials obtained from $x^{d}$ by means of infinitesimal perturbation of the coefficient $p_{d-j}$. Sum these bounds for $j=1,\dots,d$ and obtain
@@
One must access at least $0.5(d+1)db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(d+1)db$ bit operations to approximate even its single zero within $2^{-b}$ if no restriction is imposed on the number of roots in an input complex domain $\mathbb{D}$ and its $\beta$-neighborhood.
The lower bounds of this subsection imply that the algorithm of [78, 84] and its extension supporting Thm. 22.1 solve Problems 0 for $m=d$ and F by using optimal number of bit operations up to polylogarithmic factors.
Let us extend this study to Problem 0 for $m<d$. In this case isolate from one another the two sets made up of $m$ and $d-m$ roots.
Observation 22.3.
Let $p(x)=(x^{m}+p_{d-j}x^{m-j})(x-z_{1})^{d-m}$ for an integer $j$, $1<j<m$, complex $p_{d-j}$ such that $|p_{d-j}|=1/2^{bj}$, and sufficiently large $|z_{1}|$. Then the polynomial $p(x)$ has the $j$ zeros $\zeta^{g}2^{-b}$ for $g=0,1,\dots,j-1$, which vanish in the transition to $p(x)=x^{m}(x-z_{1})^{d-m}$.
Now we extend Cor. 22.2 as follows.
Corollary 22.3.
For an integer $m$, $1\leq m\leq d$, one must access at least $0.5(m+1)mb+db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(m+1)mb+0.5db$ bit operations to approximate within $2^{-b}$ even its single zero lying in the $\beta$-neighborhood of a complex domain $\mathbb{D}$ for a fixed constant $\beta>0$ if this neighborhood contains at most $m$ roots.
Proof.
Obtain the lower bounds $0.5(m+1)mb$ and $0.25(m+1)mb$ as for Cor. 22.2. Then notice that a $\delta$-perturbation of any trailing coefficient $p_{i}$ of $p(x)$ for $i\leq d-m$ can move the roots lying in the $\beta$-neighborhood of $\mathbb{D}$ by order of $\delta$. ∎
Our Main Theorem 3 implies that these lower bounds are near-optimal for $m$ of order $d$.
22.3 Linking Factorization and Root-finding: proof of Thm. 22.2
Next we prove Thm. 22.2, by applying the results of [106, Cor. 4.3 and Sec. 19].
@@
Based on [106, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
@@
[108, Thm. 2.7], which we reproduce in Sec. 23.4 as Thm. 23.3, shows an alternative way to relaxing both assumptions that $r_{1}\leq 1$ and $p(x)$ is monic.
Schönhage in [106, Sec. 19] has greatly strengthened the estimate of Thm. 22.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [106, Eqn. (19.8)]) the following result.
@@
By extending these estimates we strengthen Thm. 22.4 as follows.
@@
Under bounding cluster sizes of Def. 22.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$. Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{0}(b,d)\leq\mathbb{B}_{F}(\tilde{O}(b+d),d)$.
For converse reduction of Problem F to Problem 0 recall [106, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
@@
Now consider the products $p=gh$ for $g=x-z_{j}$, for $j=1,\dots,d$. For a fixed $j$, an error of at most $\epsilon/2^{d}$ for the root $z_{j}$ contributes at most $\epsilon||h||_{1}$ to overall factorization error of Problem F. Therefore, $||h||_{1}\leq 2^{d-1}||p||_{1}/||x-z_{j}||_{1}\leq 2^{d-1}||p||_{1}$ by virtue of Thm. 22.5. Hence the approximation errors of all the $d$ roots together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
@@
If $b=O(b-d-\log(d))$, then Problem F for an error bound $1/2^{b}$ can be reduced at a dominated bit operation cost to Problem 0 for an error bound $1/2^{\widehat{b}}$ and a sufficiently large $\widehat{b}=O(b+d+\log(d))$; hence and so $\mathbb{B}_{F}(b,d)=O(\mathbb{B}_{0}(b,d))$.
It follows that the complexity estimate of Thm. 1.3 can be also applied to the solution of Problem F of factorization provided that $b=O(b-d-\log(d))$. Then again, the bit operation complexity of this solution of Problem F is optimal (up to polylogarithmic factors) under the random root model provided that $d=\tilde{O}(b)$ (cf. Observation 22.1).
23 Recursive Factorization, its bit complexity, and Root Isolation
@@ 23.1 Auxiliary norm bounds
[108, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma 23.3. From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta_{q}$ of (3.4), one can see some limitations on strengthening this bound.
@@ 23.2 Errors and complexity
Next, by following [106, Sec. 5], we estimate the norm of the residual polynomial
@@
Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (23.3)-(23.5) and Lemmas 23.1 and 23.3 and obtain (23.6). ∎
Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (23.6) turns into (23.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (23.4), apply Thm. 23.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (23.1) with
@@
The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (23.7) ensures (23.5), and then (23.6) follows by virtue of Thm. 23.1.
@@
23.3 Overall complexity of Recursive Factorization
The overall complexity of recursive factorization is dominated by the sum of the bounds on the complexity of all splittings into pairs of factors. The algorithm of [84] adopts recursive factorization of [106] but decreases the overall complexity bound by roughly a factor of $d$. This is due to additional techniques, which ensure that in every recursive splitting of a polynomial into the product of two factors the ratio of their degrees is at most 11. Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [84], and so the overall complexity of root-finding is in $\tilde{O}(db)$ – proportional to that of the first splitting. In contrast, the degrees of a factor in any splitting in [106] can be low, e.g., 1, and so there can be up to $d-1$ splittings overall, each of the first $0.5d$ of them being almost as expensive as the very first one.
@@ 23.4 From factors to roots
By virtue of Thm. 23.3 for $b^{\prime}=O(bd)$ we can bound the bit operation complexity of the solution of Problem 0 by increasing the estimate for the complexity of factorization in Sec. 23.3 by a factor of $d$.
@@
The study in this section covers the solution of Problem 1 for the worst case input polynomial $p$. If its roots are isolated from each other, then the upper estimates of Thm. 23.3 and Cor. 23.2 can be decreased by a factor of $d$ (cf. [84]).
@@ 23.5 Polynomial Root Isolation
Based on his results recalled in this section, Schönhage in [106, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
Problem ISOL: Polynomial Root Isolation. Given a polynomial $p$ of (1.1) that has integer coefficients and is square-free, that is, has only simple roots, compute $d$ disjoint discs on the complex plane, each containing exactly one root.
@@
Problem 0 can be viewed as the Problem of Polynomial Root Cluster Isolation. It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem 0 as well as the Algorithms of [78, 84] solve Problem ISOL within a record complexity bounds.
In [106, Sec. 20] Schönhage reduces Problem ISOL to factorization Problem F. Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [60, 61, 62]). Schönhage has estimated that the solution of Problem F (e.g., presented in [106, 52, 76, 77, 84]) implies isolation provided that
@@ 24.1 Contents
In the next subsection we recall Newton’s, Weierstrass’s, and Ehrlich’s iterations and comment on their convergence and acceleration by means of fast multipoint evaluation, which is dramatic but involves the coefficients of $p$. In Sec. 24.3 we cover more limited acceleration by means of incorporation of FMM, which can be applied to a black box polynomial. In Secs. 24.5 and 25 we cover initialization of Newton’s, Weierstrass’s, and Ehrlich’s iterations, required for their fast global empirical convergence. For Newton’s iterations we only need to compute a disc that contains all zeros of a polynomial $p$ (cf. [55, 43, 115, 102, 113]). This is given by the well-known estimates via the coefficients, but we present simple novel initialization recipes in the case of a black box polynomial $p$. The customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations but is still fast for polynomials represented with their coefficients, as we recall in Sec. 25, where we also discuss applications of this initialization in particular to fast real and nearly real root-finding. It is still a challenge to extend such initialization to the case of a black box polynomial that would be faster than our radii algorithm of Sec. 16.
@@ 24.2 Definitions
Consider Newton’s, Weierstrass’s, and Ehrlich’s iterations [55, 43, 102, 115, 119, 23, 51, 28, 2, 15]):
@@
Weierstrass’s and Ehrlich’s iterations begin at $d$ points for a $d$th degree polynomial and approximate a subset of its $d$ roots about as fast and almost as slow as all its $d$ roots.
Near the roots Newton’s and Weierstrass’s iterations converge with quadratic rate, Ehrlich’s iterations with cubic rate [116]. Empirically, however, all of them consistently converge very fast globally, right from the start, if they are initialized properly. [^30] Formal support for fast global convergence is only known for for Newton’s iterations empirically they still run faster than according to the known estimates (see [102]). The known fast initialization algorithms involve the coefficients of $p$ (see [9, 15]).
All expressions in (24.3) for the functions $N(x)$, $W(x)$, and $E_{j}(x)$ can be rewritten as ratios $\frac{u(x)}{v(x)}$ for pairs of polynomials $u(x)$ and $v(x)$, whose coefficients are given or readily available at a low computational cost if the polynomial $p(x)$ is given with its coefficients (see [83, Problems 3.6.1 and 3.6.3]). Then multipoint evaluation of such a ratio amounts to multipoint evaluation of $u(x)$ and $v(x)$ and division of the computed values, and one can accelerate Newton’s, Weierstrass’s, and Ehrlich’s iterations by a factor $d/\log^{2}(d)$ by means of incorporation of the algorithms of [64, 44].